How many seconds would it take for a 4.50kw motor to raise a 175.kg boiler to a height of 15.0m?

1
Expert's answer
2022-02-20T15:45:06-0500

Given:

"P=4.50\:\rm kW"

"m=175\:\rm kg"

"h=15.0\:\rm m"


The required time

"t=\frac{W}{P}=\frac{mgh}{P}=\frac{175*9.8*15.0}{4500}=5.72\:\rm s"